You have been managing a $5 million portfolio that has a beta of 1.05 and a required rate of return of 6.675%. The current risk-free rate is 3%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 0.85, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

6.675=3+1.05*(market rate-3)

(6.675-3)=1.05*(market rate-3)

market rate=(6.675-3)/1.05+3

=6.5%

Hence required rate for $500,000=3+0.85*(6.5-3)

=5.975%

Required rate for portfolio=Respective return*Respective weight

=(5,000,000/5,500,000*6.675)+(500,000/5,500,000*5.975)

=6.61%(Approx).